# PowerInfer: Fast Large Language Model Serving with a Consumer-grade GPU

### v1

@@ 6.3 Modeling of Neuron Placement
Based on the neuron impact metric, PowerInfer utilizes a solver to optimize the total impacts of all neurons in the GPU. This cumulative impact is formulated as the objective function, as defined in Equation 2. This function is then input into an integer linear programming framework to identify a specific solution that maximizes the function. The binary variable $a_{in}$, defined in Equation 3 indicates whether the neuron $n$ is placed on processing unit $i$.
$Maximize\quad t_{i}=\sum_{e\in\mathbb{N}}a_{ie}*v_{e}\forall i\in\{GPU\}$ (2)
$\sum_{i\in\mathbb{U}}a_{in}=1\quad\forall n\in\mathbb{N}$ (3)
@@
When maximizing the objective function, the solver also needs to consider two sets of constraints associated with the communication overhead between processing units and their hardware capabilities.
6.3.1 Communication Constraint
@@
The number of neurons preloaded onto the GPU is limited by the communication overhead within a layer, a constraint dictated by hardware PCIe bandwidth limitations. If too few neurons are preloaded, this overhead negates the computational benefits offered by the GPU. As a result, the solver must identify a minimum number of neurons to allocate to the GPU for processing. This ensures that neuron processing on the GPU, including synchronization time, is more efficient than CPU processing, as specified in Inequality 4. In this inequality, $C_{l}$ is the minimum count of neurons that must be assigned to the GPU for layer $l$.
@@
When solving Inequality 4, it is essential to define both the computation time for an individual neuron in layer $l$ and the intra-layer communication overhead, $T_{sync}$. In LLM inference, especially with smaller batch sizes, the process is primarily limited by memory bandwidth. Therefore, the computation time for a neuron approximately equals the time needed to access all of its weights once, as indicated in Equation 5. With smaller batch sizes, the extent of intra-layer data transfer tends to be consistent across layers, leading to a uniform synchronization cost. Consequently, we describe $T_{sync}$ as the profiled overhead for a single instance of intra-layer communication.
$C_{l}\cdot T_{l}^{GPU}+T_{sync}\leq C_{l}\cdot T_{l}^{CPU}\forall l\in\mathbb{L}$ (4)
$T_{i}^{j}=M_{i}/Bandwidth_{j}\quad\forall j\in\mathbb{D},\forall i\in\mathbb{L}$ (5)
6.3.2 Memory Constraint
@@
To model this, we introduce an auxiliary binary variable, $y_{l}$, which can be either 1 or 0. This variable determines whether any neurons are assigned to the GPU for layer $l$. For computational convenience, a sufficiently large number $K$ is also introduced. Inequalities 7 and 8 are formulated to model this constraint. When $y_{l}$ is 1, indicating neuron placement on the GPU for this layer, and given that $K$ is adequately large, these two inequalities effectively become $y_{l}\leq\sum_{e\in N_{l}}a_{ie}\leq K$. Conversely, if $y_{l}$ is set to 0, signifying no neuron placement on the GPU for layer $l$, the inequalities reduce to $\sum_{e\in N_{l}}a_{ie}=0$.
$\sum_{n\in N}a_{jn}\cdot M_{n}<MCap_{j}\quad\forall j\in\mathbb{U}$ (6)
$\sum_{e\in N_{l}}a_{ie}\geq C_{l}\cdot y_{l}\quad\forall l\in\mathbb{L},\forall i\in\{GPU\}$ (7)
$\sum_{e\in N_{l}}a_{ie}\leq K\cdot y_{l}\quad\forall l\in\mathbb{L},\forall i\in\{GPU\}$ (8)
6.3.3 ILP Optimization
@@
Subsequently, the solver utilizes Integer Linear Programming (ILP) to optimize the objective function, conforming to all the constraints from Equation/Inequality 3 to 8. Given that ILP problems are inherently NP-complete, directly solving them for an LLM with hundreds of billions of parameters poses a considerable computational challenge. To expedite the process and achieve an approximate solution, the primary strategy involves aggregating neurons within each layer into batches for collective placement analysis. Specifically, the solver groups 64 neurons with similar impacts from a layer into a single batch. This batching strategy dramatically reduces the total neuron count, N, from several millions to roughly tens of thousands, thereby significantly decreasing the time to solve the ILP problem to approximately 10 seconds.
7 Implementation
@@
The online inference engine of PowerInfer has been implemented by incorporating an additional 4,200 lines of C++ and CUDA code into llama.cpp [14], a state-of-the-art open-source LLM inference framework designed for PCs. The extensions made by PowerInfer include modifications to the model loader for distributing an LLM across GPU and CPU, following the guidance from the offline solver’s outputs. We have also optimized the inference engine for GPU-CPU hybrid execution and introduced 10 neuron-aware operators for both processing units. All other components and functionalities of llama.cpp remains unchanged. For instance, the KV cache continues to reside in CPU memory, allowing more GPU memory for hot-activated neurons, as its access has minimal impact on inference latency, particularly in small batch sizes. Furthermore, around 400 lines of Python code were added to the transformers framework [44], enabling it to function as an offline profiler and solver for PowerInfer.
@@
The current implementation of PowerInfer supports a range of mainstream LLM families with varying parameter sizes, including the OPT [49] family (from 7B to 175B parameters), the LLaMA [42] family (7B to 70B), and Falcon-40B [2]. For these models, PowerInfer utilizes DejaVu [21] to train online activation predictors, which has been enhanced with an adaptive training method. While training an LLM is a lengthy process, often taking several hours, it is a one-time task. The duration of this process can be significantly reduced by utilizing multiple high-end GPUs.
8 Evaluation
8.1 Experimental Setup
@@
Hardware. To demonstrate the generalization of PowerInfer across various hardware setups, experiments were conducted on two distinct PC configurations, representing both high-end and low-end hardware scenarios:
@@
  • PC-High: Equipped with an Intel i9-13900K processor (eight physical cores at 5.4GHz) and 192GB host memory (memory bandwidth of 67.2 GB/s). This configuration includes an NVIDIA RTX 4090 GPU (24G) with a memory bandwidth of 1TB/s and operates with a PCIe 4.0 interface (64GB/s bandwidth).
@@
  • PC-Low: Features an Intel i7-12700K processor (eight physical cores at 4.9GHz), coupled with 64GB of host memory (memory bandwidth 38.4 GB/s). It also includes an NVIDIA RTX 2080Ti GPU (11G) with a memory bandwidth of 616GB/s and utilizes PCIe 3.0 interface (32GB/s bandwidth).
Models. We use a range of OPT [49] models with parameters from 6.7B to 175B, as well as Falcon(ReLU)-40B [38] and LLaMA(ReGLU)-70B [39] models. Notably, the 175B parameter model is comparable in size to the GPT-3 model [5]. For our experiments, all models in our experiments use FP16 and INT4 quantized parameters, with intermediate activations in FP32, consistent with recent LLM research practices [47, 12].
@@
Workloads. The workloads for our experiments are derived from ChatGPT prompts [28] and Alpaca [41] datasets, covering a wide spectrum of language model uses. These datasets consist of input and output texts typical of real LLM services. ChatGPT prompts include user interactions with ChatGPT [31], and Alpaca features instruction sets generated by GPT3.5 through self-instruction.
[…]
Key Metrics. As we focus on low latency setting, our primary evaluation metric is end-to-end generation speed, quantified as the average number of tokens generated per second (tokens/s). It is calculated by dividing the total count of generated tokens by the end-to-end response time, offering a precise measure of the response generation process’s efficiency.
8.2 End-to-End Performance
@@
Figure 10: Speedup of various models on PC-High in FP16 format. The X axis indicates the output length. The Y axis represents the speedup compared with llama.cpp. The number above each bar indicates the end-to-end generation speed (tokens/s). The first row of the figure is configured with an input length of around 64, and the second row with an input length of approximately 128.
Figure 11: Speedup of various models on PC-Low in FP16 format. The X axis indicates the output length. The Y axis represents the speedup compared with llama.cpp. The number above each bar indicates the end-to-end generation speed (tokens/s). The first row of the figure is configured with an input length of around 64, and the second row with an input length of approximately 128.
@@
We first compare the end-to-end inference performance of PowerInfer and llama.cpp with a batch size of one, the typical setting for local deployments [6]. Given real-world dialog input/output length variability [18], we sample prompts from Alpaca and ChatGPT datasets, ranging from 8 to 128 characters. Both PowerInfer and llama.cpp generated 8, 128, and 512 tokens in response to each prompt.
@@
Figure 10 illustrates the generation speeds for various models and input-output configurations on a PC-High equipped with an NVIDIA RTX 4090. On average, PowerInfer achieves a generation speed of 8.32 tokens/s, reaching up to 16.06 tokens/s, significantly outperforming llama.cpp with average speedups of 7.23$\times$, and for Falcon-40B, up to 11.69$\times$. The performance superiority of PowerInfer becomes more pronounced as the number of output tokens increases since the generation phase plays a more significant role in the overall inference time. In this phase, a small number of neurons are activated on both CPU and GPU, leading to fewer unnecessary computations compared to llama.cpp. For example, in the case of OPT-30B, only around 20% of neurons are activated for each token generated, with the majority processed on the GPU, a benefit of PowerInfer’s neuron-aware inference engine.
@@
Figure 11 shows that on a lower-end PC (PC-Low), PowerInfer still attains considerable performance enhancement over llama.cpp, averaging a speedup of 5.01$\times$ and peaking at 7.06$\times$. However, these improvements are smaller compared to those on a higher-end PC (PC-High), primarily due to the 11GB GPU memory limitation of PC-Low. This limitation affects the number of neurons that can be allocated to the GPU, particularly for models with around 30B parameters or more, leading to a greater dependence on the CPU for processing a larger number of activated neurons.
@@
(a) PC-High
(b) PC-Low
Figure 12: Neuron load distribution on CPU and GPU during inference. The yellow block refers to llama.cpp, and blue block refers to PowerInfer.
In scenarios involving long input prompts with relatively short output lengths, which are less common [28], PowerInfer demonstrates only limited performance gains. In such situations, the prompt phase, where a substantial number of tokens are processed simultaneously, becomes a crucial factor in determining inference speed. This results in each token activating a unique set of neurons, substantially diminishing activation sparsity. As a consequence, the CPU becomes the primary bottleneck in the inference process, tasked with processing a considerable number of cold-activated neurons but constrained by its computational capabilities.
Inference with Quantization. Figure 13 illustrates that PowerInfer effectively supports LLMs that are compressed using INT4 quantization. On a high-end PC (PC-High), PowerInfer delivers responses at an average speed of 13.20 tokens/s, reaching a peak of 29.08 tokens/s. The average speedup achieved compared with llama.cpp is 2.89$\times$, with a maximum of 4.28$\times$. On a lower-end setup (PC-Low), the average speedup is 5.01$\times$, peaking at 8.00$\times$. The reduction in memory requirements due to quantization enables PowerInfer to more efficiently manage larger models. For instance, in our experiment with the OPT-175B model on PC-High, PowerInfer nearly reaches two tokens per second, surpassing llama.cpp by a factor of 2.66$\times$.
@@
Figure 13: Speedup of different models on PC-High and PC-Low in INT4 format. The X axis indicates the output length. The Y axis represents the speedup compared with llama.cpp. The number above each bar indicates the end-to-end generation speed (tokens/s). The upper row of the figure presents performance on PC-High, while the lower row details those on PC-Low.
@@
Figure 14: Batch inference speedup of Falcon-40B on PC-High. The X axis indicates the request batch size, the Y axis represents the end-to-end token generation speed (tokens/s). The number above each bar shows the speedup compared with llama.cpp.
8.3 Ablation Studies
8.3.1 Performance Breakdown

### v2

@@ 6.3. Modeling of Neuron Placement
Based on the neuron impact metric, PowerInfer utilizes a solver to optimize the total impacts of all neurons in the GPU. This cumulative impact is formulated as the objective function, as defined in Equation 2. This function is then input into an integer linear programming framework to identify a specific solution that maximizes the objective function. The binary variable $a_{in}$, defined in Equation 3 indicates whether the neuron $n$ is placed on processing unit $i$.
(2) $Maximize\quad t_{i}=\sum_{e\in\mathbb{N}}a_{ie}*v_{e}\forall i\in\{GPU\}$
(3) $\sum_{i\in\mathbb{U}}a_{in}=1\quad\forall n\in\mathbb{N}$
@@
When maximizing the objective function, the solver also needs to consider two sets of constraints. First, minimize the communication overhead between processing units. Second, ensure the GPU memory is fully utilized.
6.3.1. Communication Constraint
@@
The number of neurons preloaded onto the GPU is limited by layer communication overheads, constrained by hardware PCIe bandwidth. Preloading too few neurons negates the GPU’s computational benefits. Consequently, the solver must determine a minimum neuron allocation for optimal GPU processing efficiency, as detailed in Inequality 4. In this inequality, $C_{l}$ is the minimum count of neurons that must be assigned to the GPU for layer $l$.
@@
When solving Inequality 4, it is essential to profile both the computation time for an individual neuron and the intra-layer communication overhead, $T_{sync}$. In LLM inference, particularly with smaller batch sizes, the limiting factor is memory bandwidth. Hence, a neuron’s computation time is roughly equal to the duration required to access its weights once, as shown in Equation 5. And the extent of intra-layer data transfer tends to be consistent across layers, leading to a uniform synchronization cost. Consequently, we describe $T_{sync}$ as the profiled overhead for a single instance of intra-layer communication.
(4) $C_{l}\cdot T_{l}^{GPU}+T_{sync}\leq C_{l}\cdot T_{l}^{CPU}\forall l\in\mathbb{L}$
(5) $T_{i}^{j}=M_{i}/Bandwidth_{j}\quad\forall j\in\mathbb{U},\forall i\in\mathbb{L}$
6.3.2. Memory Constraint
@@
To model this constraint, we introduce an auxiliary binary variable, $y_{l}$, which can be either 1 or 0. This variable determines whether any neurons are assigned to the GPU for layer $l$. For computational convenience, a sufficiently large number $K$ is also introduced. Inequalities 7 and 8 are formulated to model this constraint. When $y_{l}$ is 1, indicating neuron placement on the GPU for this layer, and given that $K$ is adequately large, these two inequalities effectively become $C_{l}\leq\sum_{e\in N_{l}}a_{ie}\leq K$. Conversely, if $y_{l}$ is set to 0, signifying no neuron placement on the GPU for layer $l$, the inequalities reduce to $\sum_{e\in N_{l}}a_{ie}=0$.
(6) $\sum_{n\in N}a_{jn}\cdot M_{n}<MCap_{j}\quad\forall j\in\mathbb{U}$
(7) $\sum_{e\in N_{l}}a_{ie}\geq C_{l}\cdot y_{l}\quad\forall l\in\mathbb{L},\forall i\in\{GPU\}$
(8) $\sum_{e\in N_{l}}a_{ie}\leq K\cdot y_{l}\quad\forall l\in\mathbb{L},\forall i\in\{GPU\}$
6.3.3. ILP Optimization
@@
The solver utilizes Integer Linear Programming (ILP) to maximize the objective function, conforming to all the constraints from Equation/Inequality 3 to 8. Given that ILP problems are inherently NP-complete, directly solving them for an LLM with hundreds of billions of parameters poses a considerable computational challenge. To expedite the process and achieve an approximate solution, the primary strategy involves aggregating neurons within each layer into batches for collective placement analysis. Specifically, the solver groups 64 neurons with similar impacts from a layer into a single batch. This batching strategy dramatically reduces the total neuron count, N, from several millions to roughly tens of thousands, thereby significantly decreasing the time to solve the ILP problem to approximately 10 seconds.
7. Implementation
@@
The online inference engine of PowerInfer has been implemented by incorporating 4,200 lines of C++ and CUDA code into llama.cpp (Gerganov, 2023), a state-of-the-art open-source LLM inference framework designed for PCs. The extensions made by PowerInfer include modifications to the model loader for distributing an LLM across GPU and CPU, following the guidance from the offline solver’s outputs. We have also optimized the inference engine for GPU-CPU hybrid execution and introduced 10 neuron-aware operators for both processing units. All other components and functionalities of llama.cpp remains unchanged. For instance, the KV cache continues to reside in CPU memory, allowing more GPU memory for hot-activated neurons. Furthermore, around 400 lines of Python code were added to the transformers framework (Wolf et al., 2020), enabling it to function as an offline profiler and solver for PowerInfer.
@@
The current implementation of PowerInfer supports a range of mainstream LLM families with varying parameter sizes. For these models, PowerInfer utilized general corpora such as Wikipedia (Foundation, 2024) to train online activation predictors, Notably we deliberately avoided using any of the downstream task datasets in our predictor training process. Training predictors, though time-consuming (often several hours), is a one-time task that can be expedited with multiple GPUs.
8. Evaluation
8.1. Experimental Setup
@@
Hardware. All experiments were conducted on two distinct PC configurations, representing both high-end and low-end hardware scenarios:
@@
  • PC-High: Intel i9-13900K processor (eight 5.4GHz cores), 192GB host memory (bandwidth of 67.2 GB/s), an NVIDIA RTX 4090 GPU (24GB), and PCIe 4.0 interface (64GB/s bandwidth).
@@
  • PC-Low: Intel i7-12700K processor (eight 4.9GHz cores), 64GB host memory (bandwidth 38.4 GB/s), an NVIDIA RTX 2080Ti GPU (11GB), and PCIe 3.0 interface (32GB/s bandwidth).
Models. Table 3 shows the LLMs evaluated in this section, including average activation sparsity for various LLMs in the MLP blocks. These LLMs include OPT (Zhang et al., 2022b) models with parameters from 13B to 175B, Bamboo 7B (Song et al., 2024b), as well as Falcon(ReLU)-40B (Almazrouei et al., 2023) and LLaMA(ReGLU)-70B (SparseLLM, 2024) models. In addition to LLMs using ReLU-family activation functions, we also evaluate SiLU-family LLMs, with sparsity levels of approximately 50%. For our experiments, all models in our experiments use FP16 and INT4 quantized parameters, with intermediate activations in FP32, consistent with recent LLM research practices (Yu et al., 2022; Frantar et al., 2022).
Table 3. All LLMs used for evaluation and their average activation sparsity in the MLP blocks. The sparsity is measured using the same method in Table 1.
Model Activation Function Sparsity
Bamboo-7B dReLU 90%
OPT-7B/13B/30B/66B/175B ReLU 96%-98%
Falcon(ReLU)-40B ReLU 95%
LLaMA2(ReGLU)-7B ReGLU 70%
LLaMA2(ReGLU)-13B ReGLU 78%
LLaMA2(ReGLU)-70B ReGLU 82%
Qwen1.5-4B SwiGLU 40%
LLaMA2(SwiGLU)-13B SwiGLU 43%
Yi-34B SwiGLU 53%
@@
Workloads. The workloads for our experiments are derived from chatbot arena (Zheng et al., 2023a) ChatGPT prompts (MohamedRashad, 2023) and Alpaca (Taori et al., 2023) datasets. covering a wide spectrum of language model uses. These datasets are the most representative examples of real LLM services. ChatGPT prompts include user interactions with ChatGPT (OpenAI, 2023), and Alpaca features instruction sets generated by GPT3.5 through self-instruction.
Baseline System. We compare PowerInfer with llama.cpp (Gerganov, 2023) and SpecInfer (Miao et al., 2023), state-of-the-art local LLM inference frameworks. llama.cpp is the most widely used LLM inference framework for local scenarios. For a fair comparison, we extended llama.cpp to support the OPT model, as it does not natively do so. SpecInfer is representative speculative decoding framework, that utilizes smaller draft models to generate tokens and subsequently verifies these tokens in batches using the original model. While other alternatives like FlexGen (Sheng et al., 2023b) and DejaVu (Liu et al., 2023b) exist, they exhibit higher latency in the latency-sensitive scenarios discussed in this paper, as analyzed in §$\S$2.2.
8.2. End-to-End Performance
@@
Figure 10. Speedup of various models on PC-High in FP16 format. The X axis and Y axis indicate the output length and speedup compared with llama.cpp. The number above each bar indicates the end-to-end generation speed (tokens/s). The up and down figures represent input length of 64 and 128.
Figure 11. Speedup of various models on PC-Low in FP16 format. The up and down figures represent input length of 64 and 128, respectively.
@@
We first compare the end-to-end inference performance of PowerInfer , llama.cpp and SpecInfer with a batch size of one, the typical setting for local deployments (Cai et al., 2023). Given real-world dialog input/output length variability (Kwon et al., 2023), we sample prompts from Alpaca and ChatGPT datasets, ranging from 8 to 128 tokens and measure the generation speed.
@@
Figure 8.2 illustrates the generation speeds for various models and input-output configurations on a PC-High equipped with an NVIDIA RTX 4090. On average, PowerInfer achieves a generation speed of 8.32 tokens/s, reaching up to 16.06 tokens/s, significantly outperforming llama.cpp and SpecInfer with average speedups of 7.23$\times$ and 6.19$\times$, and for Falcon-40B, up to 11.69$\times$ compared with llama.cpp. Although SpecInfer utilizes speculative decoding, since the large model size exceeds the GPU’s capacity, there is still a significant amount of swapping during the verify phase, accounting for over 95% of the time. As a result, SpecInfer does not significantly outperform llama.cpp. The performance superiority of PowerInfer becomes more pronounced as the number of output tokens increases since the generation phase plays a more significant role in the overall inference time. In this phase, a small number of neurons are activated on both CPU and GPU, leading to fewer unnecessary computations compared to llama.cpp. For example, in the case of OPT-30B, only around 20% of neurons are activated for each token generated, with the majority processed on the GPU, a benefit of PowerInfer’s neuron-aware inference engine.
@@
Figure 8.2 shows that on a lower-end PC (PC-Low), PowerInfer still attains considerable performance enhancement over llama.cpp and SpecInfer, averaging a speedup of 4.71$\times$, 5.97$\times$ and peaking at 7.06$\times$ and 7.47$\times$. However, these improvements are smaller compared to those on a higher-end PC (PC-High), primarily due to the 11GB GPU memory limitation of PC-Low. This limitation affects the number of neurons that can be allocated to the GPU, particularly for models with around 30B parameters or more, leading to a greater dependence on the CPU for processing a larger number of activated neurons.
Figure 12. Neuron load distribution on CPU and GPU during inference. The yellow block refers to llama.cpp, and blue block refers to PowerInfer.
@@
Inference with different input length. In scenarios involving long input prompts with extremely short output lengths (less than 8 tokens), which are less common (MohamedRashad, 2023), PowerInfer demonstrates limited performance gains, ranging from 1.07$\times$ (Figure 8.2) to 4$\times$ (Figure 8.2). In such situations, the prompt phase, where a substantial number of tokens are processed simultaneously, becomes a crucial factor in determining inference speed. This results in each token activating a unique set of neurons, substantially diminishing activation sparsity. As a consequence, the CPU becomes the primary bottleneck in the inference process, tasked with processing a considerable number of cold-activated neurons but constrained by its computational capabilities.
For relatively long input sequences, PowerInfer still achieves a 3.47$\times$ to 5.69$\times$ speedup, as shown in Table 4. This is because when the input sequence is lengthy, all neurons are activated collectively by the input tokens. In this case, PowerInfer switches to dense GPU computation during the prefill stage, resulting in prefill latency comparable to llama.cpp. During the generation phase, where only one token is processed at a time, sparsity emerges in each inference step. Here, PowerInfer leverages its hybrid engine to achieve significant acceleration.
Table 4. End-to-end latency comparison for different models with 1.5K input length and 256 output length on PC-High.
LLMs llama.cpp (ms) PowerInfer (ms) Speedup
LLaMA(ReGLU)-13B-FP16 49.91 14.38 3.47$\times$
Falcon(ReLU)-40B-FP16 321.63 56.48 5.69$\times$
LLaMA(ReGLU)-70B-FP16 92.76 37.17 2.50$\times$
Inference with Quantization. Figure 13 illustrates that PowerInfer effectively supports LLMs that are compressed using INT4 quantization. We fail to run SpecInfer with INT4 quantization due to the lack of support. On a high-end PC (PC-High), PowerInfer delivers responses at an average speed of 13.20 tokens/s, reaching a peak of 29.08 tokens/s. The average speedup achieved compared with llama.cpp is 2.89$\times$, with a maximum of 4.28$\times$. On a lower-end setup (PC-Low), the average speedup is 5.01$\times$, peaking at 8.00$\times$. The reduction in memory requirements due to quantization enables PowerInfer to more efficiently manage larger models. For instance, in our experiment with the OPT-175B model on PC-High, PowerInfer nearly reaches two tokens per second, surpassing llama.cpp by a factor of 2.66$\times$.
@@
Figure 13. Speedup of different models on PC-High and PC-Low in INT4 format. The upper row of the figure presents performance on PC-High, while the lower row details those on PC-Low.
@@
Figure 14. Batch inference speedup of Falcon-40B on PC-High. The X axis indicates the request batch size, the Y axis represents the end-to-end token generation speed (tokens/s). The number above each bar shows the speedup compared with llama.cpp.
8.3. Ablation Studies
8.3.1. Performance Breakdown